Grade 11Mechanics

A point moves in x-y plane according to law x=4sin6t and y=4(1-cos6t).the distance traversed by particle in 4 seconds is where x y are in meter

adesh Singalwar
10 Years agoGrade 11

1 Answer

Abhishek
7 Years ago
By given information it form a circle centered at (0,4) with radius 4 at t=0  position of particle is(0,0) angular position=6*4 =24 therefore s= r.angular position =4*24=96